# Grounding Video Models to Actions through Goal Conditioned Exploration

### v1

@@ 2 Related Work
A large body of recent work has explored how video models can be used in decision making (Yang et al., 2024d; McCarthy et al., 2024). Prior work has explored how video models can act as reward functions (Escontrela et al., 2024; Huang et al., 2023; Chen et al., 2021), representation learning (Seo et al., 2022; Wu et al., 2023; 2024; Yang et al., 2024b), policies (Ajay et al., 2024; Du et al., 2024; Liang et al., 2024), dynamics models (Yang et al., 2024c; Du et al., 2024; Zhou et al., 2024b; Brooks et al., 2024; Rybkin et al., 2018; Mendonca et al., 2023), and environments (Bruce et al., 2024). Our work explores that given video generations, how we can learn a policy and infer the actions to execute in an environment without any action labels, while existing works (Baker et al., 2022; Bharadhwaj et al., 2024a; Black et al., 2023; Wen et al., 2023; Zhou et al., 2024a; Wang et al., 2024) requires domain specific action data. Most similar to our work is AVDC (Ko et al., 2023), which uses rigid object transformations computed by optical flow between video frames as an approach to extract actions from generated videos. We propose an alternative unsupervised approach to ground video models to continuous action by leveraging video-guided goal-conditioned exploration to learn a goal-conditioned policy.
@@ 3 Method
3.1 Learning Goal-Conditioned Policies through Video-Guidance and Hindsight Relabeling
@@
Video models have emerged as a powerful source of prior knowledge about the world, providing rich information about how to complete various tasks from large-scale internet data. We leverage the knowledge contained in these models to help guide our exploration in an environment to solve new tasks. To this end, we propose a novel method that uses a pre-trained video model $f_{\theta}(x_{\text{start}},c)$ to guide the exploration and shrink the search space only to task-relevant states ($x_{\text{start}}$ denote the initial observation and $c$ denote the corresponding task description). This concurrently benefits sides: the goal-conditioned policy obtains task-relevant goals so as can perform efficient exploration centered around the task-relevant state space; on the other hand, the underlying information from the video model is extracted to end effectors and enables effective decision-making and control in embodied agents. Specifically, during exploration, we first leverage the video model to generate a sequence of images based on the given image observation $x_{\text{start}}$ and language task description $c$,
@@ 3.2 Periodic Random Action Bootstrapping
where $n_{r}$ denotes the number of random action episodes, $t^{r}$ denotes the length of a random action episode, $a_{\text{low}}$ and $a_{\text{high}}$ are the action limits. This proposed bootstrapping method can enhance the exploration in two ways: the initial random actions serve as the basic world dynamics information which enables the policy to reach the vicinity of goal states specified by video frames, ensuring effective video-guided exploration; the periodic extra random exploration can further expand the agent’s discovered state space and stabilize policy training. Please see Section 4.3 for ablation studies.
Algorithm 1 Grounding Video Models to Actions
1:Require: a frozen video diffusion model $f_{\theta}(x_{\text{start}},c)$, a goal-conditioned policy to train $\pi(a|x_{\text{start}},x_{\text{goal}})$, a replay buffer $R$
2:Hyperparameters: horizon of policy $\pi$ $h$, training iteration $N$, number of initial random action episodes $n_{r}$, video-guided exploration frequency $q_{v}$
@@
5:     if  $i\mod q_{v}$ == 0 then
@@
10:     $(x_{i:i+h},a_{i:i+h})$ = sample a consecutive sequence of image-action pairs from replay buffer $R$
11:     $a^{\text{pred}}=\pi(a|x_{i},x_{i+h})$
12:     $\mathcal{L}=\text{MSE}(a^{\text{pred}},a_{i:i+h})$
13:end for
14:return goal-conditioned policy $\pi$
@@ 3.3 Chunk Level Action Prediction and Exploration
Figure 2: Environment Demonstrations. We evaluate our method on three robotic manipulation environments: Libero, Meta-World, Calvin, and one visual navigation environment: iThor, with a total of 30 tasks.
@@ 4.1 Manipulation
In this section, we aim to evaluate the goal-conditioned policy learned by the proposed unsupervised exploration in tabletop manipulation environments with continuous action space. To better understand the capability of the method, we investigate multi-task policy learning in Libero and Calvin, and single-task policy learning in Meta-world. Note that methods highlighted by an asterisk require ground-truth action labels to train, while our method only requires image observations for training the video model with the potential of zero-shot generalization if trained on internet-scale data.
@@
Table 2: Quantitative Results of 6 tasks in Meta-World. Methods marked with an asterisk ‘*’ require ground-truth action demonstrations to train. AVDC uses the segmentation mask of the target object to compute actions. Our method learns the policy by video-guided self-exploration in the environment without any external supervision.
@@
Figure 7: Qualitative Results of Navigation to Bread in iTHOR environment. The start position of the agent is randomized in test time. Only a subset of the predicted video frames are shown due to the space limit. In navigation tasks with a moving camera, the video model can still generate realistic frames that respect the the actual scene layout. Our goal-conditioned policy is able to follow the video frames and reach the bread shown in the middle of the frame. However, BC turns to the other way, and AVDC cannot correctly infer the required actions and miss the target.
@@ 4.2 Visual Navigation
iTHOR (Kolve et al., 2017) is a room-level vision-based simulated environment, where agents can navigate in the scenes and interact with objects. We adopt the iTHOR visual object navigation benchmark, where agents are required to navigate to the specific type of objects given by a nature language input. The action space consists of four actions: Move Ahead, Turn Left, Turn Right, and Done. We incorporate 4 different scene types: Kitchen, Living Room, Bedroom, and Bathroom, with 3 targets in each scene. Following (Ko et al., 2023), the video model is trained on 20 demonstration image sequences per task. BC and GCBC are also trained on the same demonstrations while with access to corresponding actions. Since the navigation actions are discrete and episode lengths (typically $<20$) are much shorter than robot arm manipulation tasks, we find that predicting a single action suffices. Our method uses the same ResNet+MLP model architecture as BC-based baselines.
@@ 5 Discussion
Our approach has several limitations. First, since the approach relies on goal-conditioned random exploration and without access to action labels, for tasks that require very precise manipulation (i.e. stacking a block tower in millimeter precision), our random exploration procedure may not find the precise set of actions. In such settings, having a random exploration primitive (i.e. stacking a block on top of another) on which we do goal-conditioned policy learning may help us find such precise actions. In addition, purely random exploration in the physical world might pose additional requirements for the workbench, as sampled actions or the initial actions outputted by the learned goal-conditioned policy may cause undesired contacts between the robot and the external environment. This can be partially mitigated by integrating hard-coded safety constraints and is also a further direction of future work. Finally, our method usually requires dozens of video-guided rollouts to obtain a competent policy for a task. While resetting an environment is easy in simulation, this might demand additional work in real-world experiments. Combining our approach with a policy that can reset environment states would be an interesting direction for the future.
@@ Appendix A Additional Quantitative Results
In this section, we provide additional quantitative experiment results. In Section A.1, we compare our methods with BC trained on additional demonstration data. Next, in Section A.2, we investigate the effect of training the video model with different amounts of demonstration data. In Section A.3, we experiment our proposed method with video models of different horizons. Then, in Section A.4, we study how varying the video exploration frequency $q_{v}$ affects exploration performance. In Section A.5, we conduct ablation studies of using goal-conditioned policies with different horizons. Following this, in Section A.6, we evaluate the policy after different numbers of training iterations. In Section A.7, we compare our method with a reinforcement learning baseline with zero-shot reward. Finally, in Section A.8, we conduct ablation studies on the task success rate over the number of video-guided exploration rollouts.
@@ D.2 Training Details
In training, we randomly sample sequences of image-action pairs of horizon $h$ from the replay buffer. For Diffusion Policy setup, we mainly follow the hyperparameters suggested in the original paper. We provide detailed hyperparameters for training our model in Table 12 and 13. We do not apply any hyperparameter search or learning rate scheduler. For Libero environment, the training time of our model is approximately 36 hours for a full 200k training steps. However, the performance can gradually saturate with much fewer steps. For the Meta-World and iThor environments, the training time of our model is approximately one day; for Calvin, the training time of our model is around 15 hours. For these environments, we also observe that the performance can gradually saturate within less training time.
@@
D.3 Implementation of Baselines

### v2

@@ 2 Related Work
A large body of recent work has explored how video models can be used in decision making (Yang et al., 2024d; McCarthy et al., 2024). Prior work has explored how video models can act as reward functions (Escontrela et al., 2024; Huang et al., 2023; Chen et al., 2021; Luo et al., 2024), representation learning (Seo et al., 2022; Wu et al., 2023; 2024; Yang et al., 2024b), policies (Ajay et al., 2024; Du et al., 2024; Liang et al., 2024), dynamics models (Yang et al., 2024c; Du et al., 2024; Zhou et al., 2024b; Brooks et al., 2024; Rybkin et al., 2018; Mendonca et al., 2023), and environments (Bruce et al., 2024). Our work explores that given video generations, how we can learn a policy and infer the actions to execute in an environment without any action labels, while existing work (Baker et al., 2022; Bharadhwaj et al., 2024a; Black et al., 2023; Wen et al., 2023; Zhou et al., 2024a; Wang et al., 2024) requires domain specific action data. Most similar to our work is AVDC (Ko et al., 2023), which uses rigid object transformations computed by optical flow between video frames as an approach to extract actions from generated videos. We propose an alternative unsupervised approach to ground video models to continuous action by leveraging video-guided goal-conditioned exploration to learn a goal-conditioned policy.
@@ 3 Method
3.1 Learning Goal-Conditioned Policies through Hindsight Relabeling
@@
Video models have emerged as a powerful source of prior knowledge about the world, providing rich information about how to complete various tasks from large-scale internet data. We leverage the knowledge contained in these models to help guide our exploration in an environment to solve new tasks. To this end, we propose a novel method that uses a pre-trained video model $f_{\theta}(x_{\text{start}},c)$ to guide the exploration and shrink the search space only to task-relevant states ($x_{\text{start}}$ denote the initial observation and $c$ denote the corresponding task description). This concurrently benefits both sides: the goal-conditioned policy obtains task-relevant goals so as can perform efficient exploration centered around the task-relevant state space; on the other hand, the underlying information from the video model is extracted to end effectors and enables effective decision-making and control in embodied agents. Specifically, during exploration, we first leverage the video model to generate a sequence of images based on the given image observation $x_{\text{start}}$ and language task description $c$,
@@ 3.2 Periodic Random Action Bootstrapping
where $n_{r}$ denotes the number of random action episodes, $t^{r}$ denotes the length of a random action episode, $a_{\text{low}}$ and $a_{\text{high}}$ are the action limits. This proposed bootstrapping method can enhance the exploration in two ways: the initial random actions serve as the basic world dynamics information which enables the policy to reach the vicinity of goal states specified by video frames, ensuring effective video-guided exploration; the periodic extra random exploration can further expand the agent’s discovered state space and stabilize policy training. Please see Section 4.3 for ablation studies and Appendix D.3 for implementation details.
Algorithm 1 Grounding Video Model to Actions
1:Require: a frozen video diffusion model $f_{\theta}(x_{\text{start}},c)$, a goal conditioned policy to train $\pi(a|x_{\text{start}},x_{\text{goal}})$, a replay buffer $R$
2:Hyperparameters: horizon of policy $\pi$ $h$, training iteration $N$, number of initial / additional random action episodes $n_{r}$ / $n_{r}^{\prime}$, video-guided / random action exploration frequency $q_{v}$ / $q_{r}$
@@
5:     if  $i\mod q_{v}$ == 0 then  # conduct video-guided exploration with frequency $q_{v}$
@@
10:     if  $i\mod q_{r}$ == 0 then  # conduct periodic random action bootstrapping with frequency $q_{r}$
11:         Sample additional $n_{r}^{\prime}$ episodes of random actions and add to replay buffer $R$
12:     end if
13:     # train the policy with the data sampled from the replay buffer
14:     $(x_{i:i+h},a_{i:i+h})$ = sample a consecutive sequence of image-action pairs from replay buffer $R$
15:     $a^{\text{pred}}=\pi(a|x_{i},x_{i+h})$
16:     $\mathcal{L}=\text{MSE}(a^{\text{pred}},a_{i:i+h})$
17:end for
18:return goal-conditioned policy $\pi$
@@ 3.3 Chunk Level Action Prediction and Exploration
Figure 2: Environment Demonstrations. We evaluate our method on three robotic manipulation environments: Libero, Meta-World, Calvin, and one visual navigation environment: iThor, with a total of 30 tasks. Images in (a) and (c) denote the goal object states of a subset of tasks. Images in (b) are randomly sampled start observations of a subset of tasks. In (d), we show the layout for each scene from the agent’s view.
@@ 4.1 Manipulation
In this section, we aim to evaluate the goal-conditioned policy learned by the proposed unsupervised exploration in tabletop manipulation environments with continuous action space. To better understand the capability of the method, we investigate multi-task policy learning in Libero and Calvin, and single-task policy learning in Meta-world. Note that methods highlighted by an asterisk require ground-truth action labels to train, whereas our method only requires image demonstration sequences to train the video model and self-supervised training for the policy model.
@@
Table 2: Quantitative Results of 6 tasks in Meta-World. Methods marked with ‘*’ requires ground-truth action demonstrations to train. AVDC uses the segmentation mask of the target object to compute actions. Our method learns the policy by video-guided self-exploration in the environment without any external supervision.
@@
Figure 7: Qualitative Results of Navigation to Bread in iTHOR environment. Only a subset of the predicted video frames are shown due to the space limit. In navigation tasks with a moving camera, the video model can still generate realistic frames that respect the actual scene layout. Our goal-conditioned policy is able to follow the video frames and reach the bread shown in the middle of the frame. However, BC turns to the other way, and AVDC cannot correctly infer the required actions and miss the target.
@@ 4.2 Visual Navigation
iTHOR (Kolve et al., 2017) is a room-level vision-based simulated environment, where agents can navigate in the scenes and interact with objects. We adopt the iTHOR visual object navigation benchmark, where agents are required to navigate to the specific type of objects given by a natural language input. The action space consists of four actions: Move Ahead, Turn Left, Turn Right, and Done. We incorporate 4 different scene types: Kitchen, Living Room, Bedroom, and Bathroom, with 3 targets in each scene. Following (Ko et al., 2023), the video model is trained on 20 demonstration image sequences per task. BC and GCBC are also trained on the same demonstrations while with access to corresponding actions. Since the navigation actions are discrete and episode lengths (typically $<20$) are much shorter than robot arm manipulation tasks, we find that predicting a single action suffices. Our method uses the same ResNet+MLP model architecture as BC-based baselines.
@@ 5 Discussion
Our approach has several limitations. First, since the approach relies on goal-conditioned random exploration, for tasks that require very precise manipulation (i.e. stacking a block tower in millimeter precision), our random exploration procedure may not find the precise set of actions. In such settings, having a random exploration primitive (i.e. stacking a block on top of another) on which we do goal-conditioned policy learning may help us find such precise actions. In addition, purely random exploration in the physical world might pose additional requirements for the workbench, as sampled actions or the initial actions outputted by the learned goal-conditioned policy may cause undesired contacts between the robot and the external environment. This can be partially mitigated by integrating hard-coded safety constraints and is also a further direction of future work. Finally, our method requires dozens of video-guided rollouts to obtain a competent policy for a task. While resetting an environment to obtain these rollouts is easy in simulation, exploring how to autonomously reset and explore in the real world is a direction of future research.
@@ Appendix A Additional Quantitative Results
In this section, we provide additional quantitative experiment results. In Section A.1, we compare our methods with BC trained on additional demonstration data. Next, in Section A.2, we investigate the effect of training the video model with different amounts of demonstration data. In Section A.3, we experiment our proposed method with video models of different horizons. Then, in Section A.4, we study how varying the video exploration frequency $q_{v}$ affects exploration performance. In Section A.5, we conduct ablation studies of using goal-conditioned policies with different horizons. Following this, in Section A.6, we evaluate the policy after different numbers of training iterations. In Section A.7, we compare our method with a reinforcement learning baseline with zero-shot reward. In Section A.8, we conduct ablation studies on the task success rate over the number of video-guided exploration rollouts. Finally, in Section A.9, we study the robustness of the resulting goal-conditioned policy learned by exploration to video models of different sampling rates.
@@ A.7 Comparison to RL with Zero-Shot Reward
A.9 Policy Robustness to Video Models of Different Sampling Rates
In this section, we study the robustness of the resulting goal-conditioned policy learned by exploration to video models of different sampling rates. Specifically, once the policy is trained, we replace the training-time video model with video models of different sampling rates while remain using the same policy.
Following the video model design in Ko et al. (2023), we use the Number of Frames to Goal to denote the sampling rate of the video model. For example, Number of Frames to Goal = 9 indicates that the video model is designed to uniformly generate 9 frames from the initial observation to the goal state, that is, compared to a 7-frame video model, the sampling rate is denser (i.e., the temporal distance between two adjacent frames is smaller).
We present the policy success rate in Table A.9. In this table, the policy is trained with a 7-frame video model (marked with an asterisk) while the number of frames to goal of the test-time video model varies from 5 to 9. The policy consistently achieves high performance across video models with different sampling rates, demonstrating its robustness to variations in the sampling rate during testing. This is probably because of the robust policy training enabled by the combination of random action bootstrapping and video-guided exploration.
We observe that the 6-frame video model attains the highest success rate, even surpassing the training-time 7-frame model. This improvement is likely due to the better quality of the synthesized video frames, as modeling a greater or denser number of frames increases complexity while we keep the video model size unchanged in this experiment.
{tabu}
c c c c c c # Frames to Goal put-red-mug-left put-red-mug-right put-white-mug-left put-Y/W-mug-right Overall
5 32.8±$\pm\mathtt{12.0}$ 46.4±$\pm\mathtt{10.3}$ 66.4±$\pm\mathtt{5.4}$ 41.6±$\pm\mathtt{4.8}$ 46.8±$\pm\mathtt{8.1}$
6 38.4±$\pm\mathtt{4.8}$ 48.0±$\pm\mathtt{8.8}$ 68.8±$\pm\mathtt{4.7}$ 36.0±$\pm\mathtt{6.7}$ 47.8±$\pm\mathtt{6.2}$
  7* 38.4±$\pm\mathtt{15.3}$ 40.8±$\pm\mathtt{7.8}$ 51.2±$\pm\mathtt{3.9}$ 38.4±$\pm\mathtt{8.6}$ 42.2±$\pm\mathtt{8.9}$
8 39.2±$\pm\mathtt{6.4}$ 51.2±$\pm\mathtt{13.7}$ 65.6±$\pm\mathtt{4.8}$ 25.6±$\pm\mathtt{5.4}$ 45.4±$\pm\mathtt{7.6}$
9 32.8±$\pm\mathtt{5.3}$ 44.0±$\pm\mathtt{8.8}$ 36.8±$\pm\mathtt{9.9}$ 28.8±$\pm\mathtt{5.9}$ 35.6±$\pm\mathtt{7.5}$
{tabu}
to c c c c c c
# Frames to Goal put-choc-left put-choc-right put-red-mug-plate put-white-mug-plate Overall
5 62.4±$\pm\mathtt{4.1}$ 76.8±$\pm\mathtt{11.7}$ 66.4±$\pm\mathtt{4.1}$ 59.2±$\pm\mathtt{10.6}$ 66.2±$\pm\mathtt{7.6}$
6 64.8±$\pm\mathtt{6.9}$ 72.0±$\pm\mathtt{12.4}$ 74.4±$\pm\mathtt{9.3}$ 58.4±$\pm\mathtt{9.3}$ 67.4±$\pm\mathtt{9.5}$
  7* 70.4±$\pm\mathtt{12.8}$ 79.2±$\pm\mathtt{3.9}$ 72.8±$\pm\mathtt{6.4}$ 25.6±$\pm\mathtt{11.5}$ 62.0±$\pm\mathtt{8.7}$
8 52.0±$\pm\mathtt{12.1}$ 84.8±$\pm\mathtt{8.2}$ 68.0±$\pm\mathtt{8.4}$ 26.4±$\pm\mathtt{10.9}$ 57.8±$\pm\mathtt{9.9}$
9 59.2±$\pm\mathtt{13.9}$ 76.0±$\pm\mathtt{8.8}$ 74.4±$\pm\mathtt{9.7}$ 20.8±$\pm\mathtt{7.3}$ 57.6±$\pm\mathtt{9.9}$
Table 12: Policy Performance with Video Models of Different Sampling Rates at Test Time. All results in the table are from the same goal-conditioned policy but with video models of different sampling rates (identified by the number of intermediate frames from initial observation to the goal state). The policy uses a video model of # Frames to Goal = 7 to conduct video-guided exploration during training, but is evaluated with video models of different numbers of to-goal frames at test time. The policy performs consistently and is robust to various sampling rates at test time.
@@ D.2 Training Details
In training, we randomly sample sequences of image-action pairs of horizon $h$ from the replay buffer. For Diffusion Policy setup, we mainly follow the hyperparameters suggested in the original paper. We provide detailed hyperparameters for training our model in Table 13 and 14. We do not apply any hyperparameter search or learning rate scheduler. For Libero environment, the training time of our model is approximately 36 hours for a full 200k training steps. However, the performance can gradually saturate with much fewer steps. For the Meta-World and iThor environments, the training time of our model is approximately one day; for Calvin, the training time of our model is around 15 hours. For these environments, we also observe that the performance can gradually saturate within less training time.
@@
D.3 Random Action Bootstrapping Details
Random Action Bootstrapping. We provide detailed hyperparameters for random action bootstrapping in Table 15, where we use the same notations as in Section 3 and Algorithm 1.
In Table 15, # of Initial Episodes $n_{r}$ and # of Addition Episodes $n_{r}^{\prime}$ are with respect to one task for tabletop manipulation and with respect to one scene in iTHOR navigation. In Libero, the policy learns all 8 tasks concurrently; In Meta-World, the policy learns each task separately; In Calvin, the policy learns all 4 tasks concurrently; In iTHOR, the policy learns all 4 scenes (12 tasks) concurrently. In addition, before video-guided exploration, we first warm-start the policy by training it solely on the initial random action episodes for $N_{r}$ steps.
We use the first in, first Out convention to implement the replay buffer. The replay buffer is shared across tasks if doing multi-task exploration, thus the replay buffer size is relatively larger for multi-task setting.
Libero Meta-World Calvin iTHOR
Warm-start steps $N_{r}$ 10k 10k 10k 10k
Episode Length 120 120 120 50
Action Chunk Size $l_{c}$ 24 24 24 1
# of Initial Episodes $n_{r}$ 50 200 100 100
Periodic Frequency $q_{r}$ 500 500 500 500
# of Addition Episodes $n_{r}^{\prime}$ 2 10 4 2
Replay Buffer Size 1200 500 1200 1200
Table 15: Hyperparameters for Random Action Bootstrapping in each Environment. Same notations are used as in Section 3 and Algorithm 1. $n_{r}$ and $n_{r}^{\prime}$ represent number of random action exploration per task.
In Table 16, we provide the values of $a_{\text{low}}$ and $a_{\text{high}}$ for sampling random actions, which are introduced in Section 3.2. For visual navigation tasks, since the action space is discrete, we sample from the four possible actions {Move Ahead, Turn Left, Turn Right, Done} with uniform probability during the random action bootstrapping.
Libero Meta-World Calvin
$a_{\text{low}}$ [-1,-1,-1,-0.01,-0.01,-0.01] [-1,-1,-1] [-1,-1,-1,-1,-1,-1,-1]
$a_{\text{high}}$ [1,1,1,0.01,0.01,0.01] [1,1,1] [1,1,1,1,1,1]
Table 16: Values of $a_{\text{low}}$ and $a_{\text{high}}$ for Random Action Sampling. The gripper action is discretized to -1 and 1 (close and open) hence not included in the table. In Libero environment, a smaller value in the orientation action space is set for more efficient random action bootstrapping.
D.4 Task Success Metric
In this section, we describe the metric to evaluate whether a task is successfully completed. Specifically, the simulation environment will check the state of the target object in manipulation tasks or check the agent state in visual navigation tasks after executing an action. For manipulation tasks, if the target object is within the range of the specified goal state, the environment will return a success; for navigation tasks, if the target object is within a certain distance to and in the view of the agent and the agent executes a ’Done’ action, the environment will return a success.
During a rollout, if a success is returned, this rollout will be counted as success and terminated; otherwise, if we have finished a rollout (i.e., have sequentially executed all the synthesized video frames) and no success is returned, this episode will be counted as a failed rollout.
D.5 Implementation of Baselines